OpenStudy (linn99123):

Simplify by multiplying and adding like terms. 6(3x) + 5(2x) = @iGreen @undeadknight26

OpenStudy (linn99123):

ill brb mom calling me

OpenStudy (igreen):

Just multiply 6 and 5 into the parenthesis. \((6 \cdot 3x = ?) + (5 \cdot 2x = ?)\)

OpenStudy (studygurl14):